OpenStudy (anonymous):

Find a quadratic model for the set of values: (-2, -20), (0,-4), (4,-20). Show your work

OpenStudy (loser66):

What is the quadratic model?

OpenStudy (anonymous):

Let the quadratic model be \[y=ax^2+bx+c~ ..... (1)\] when x=0,y=-4 -4=0+0+c c=-4 \[when~x=-2,y=-20\] \[-20=4a-2b-4\] when x=4,y=-20 -20=16a+4b-4 16a+4b-4=4a-2b-4 12a=-6b -2a=b -20=4a-2b-4 -16=4a-2(-2a) 8a=-16 a=-2 b=-2*-2=4 y=-2x^2+4x-4